Define the cumulative distribution function F(t) oft by F(t) = P(W <t). Shade the region consisting of each point (x, y) in wLet X be binomial with parameters n = 10 and p = 0.5. Compute P(X = 5). Compute P(3 sX s 6). Compute P(X s 3). Submit You havThe annual income of graduates from college H in their first year of employment is normally distributed with mean of $45,000Exactly $46,000. Greater than $65,000.
Answer #1

given = 1-0.5 10-5 p(x=5) = 10es -> (120).s.cos) + 210. + costs x be binomial with parameters n=1098=65 Anubhchip) The prob1 (120) - + 210. 250... 32 16 32 . +216. 64 -) 120 252 216 + 210 (०१५ न 1024 २०२५ 1624 792 104 ११ 128 10-1 - 0 1734 -. (8369compate pex=5) 0.2460 computi p(3<356) 0.7734 compute CX =3) Lo @.1719meinwa given normally distribution with $45,000 standand clication (G) = $10,000 less than $35,500 - P(x 4835,000 (X-11 < 336Between us out and 6000 peesoonex <GOWO) - prasto-, med Gorbo!) = 45000-4000 60000-4560 << 10000 1600 << 15000 lowo 10000 apcp(z> 65000–45000 - pezs 20000 1000 - plaza) -la-pez<2) Using normal &- table 0.9772 0.0 228 ...pex765000) = 10.00 28 - less t